References

[1] Erol Akarsu, Geoffrey Fox, and Tomasz Haupt. DARP: Java-based data analysis and rapid prototyping environment for distributed high performance computations. In Fox [5]. To appear in Concurrency: Practice and Experience.

[2] Mark Baker, Bryan Carpenter, Sung Hoon Ko, and Xinying Li. mpi-Java: A Java interface to MPI. To be presented at First UK Workshop on Java for High Performance Network Computing, Europar 1998. http://www.npac.syr.edu/projects/pcrc/docs.html.

[3] Bryan Carpenter, Geoffrey Fox, Donald Leskiw, Xinying Li, Yuhong Wen, and Guansong Zhang. Language bindings for a data-parallel runtime. In Michael Gerndt and Hermann Hellwagner, editors, Third International Workshop on High-Level Parallel Programming Models and Supportive Environments. IEEE Computer Society Press, 1998.

[4] Bryan Carpenter, Guansong Zhang, Geoffrey Fox, Xiaoming Li, Xinying Li, and Yuhong Wen. Towards a Java environment for SPMD programming. To be presented at Europar ’98. http://www.npac.syr.edu/projects/pcrc/docs.html.

[5] Geoffrey C. Fox, editor. ACM 1998 Workshop on Java for High-Performance Network Computing, February 1998. To appear in Concurrency: Practice and Experience.

[6] Tomasz Haupt, Erol Akarsu, Geoffrey Fox, and Wojtek Furmanski. Web based metacomuting. http://www.npac.syr.edu/users/haupt/WebFlow/papers/FGCS.